stan idon't knowwhat youmean by the answer you gave me on lenear equations
8x-4y=16
4x=5y=22
you said multiply 2nd by 2 to get 3rd 8x+5y=22 how did you get 2 and so on i need more explanation or i wont understand

Answer by checkley75(3666)

8x-4y=16
4x+5y=22 if this is your second equation ( first= should be a +???) then
ultiply the second equation by -2 and add.
8x-4y=16
-8x-10y=-44
-------------------------------------
-14y=-28
y=-28/-14
y=2 answer. now substitute -2 for y & solve for x
8x-4*2=16
8x-8=16
8x=16+8
8x=24
x=24/8
x=3 answer.
proof
4*3+5*2=22
12+10=22
22=22